OpenStudy (sanra123):

Fay threw a basketball from the basketball court toward the hoop. The quadratic equation that models the height of the ball after t seconds is(hx)=−16t2 +20t+6.if the hoop is 10 feet high, how long is the ball in the air before the ball goes through the hoop?

OpenStudy (sanra123):

@ganeshie8

ganeshie8 (ganeshie8):

set the height function equal to 10 and solve t

OpenStudy (sanra123):

How do we solve for t

ganeshie8 (ganeshie8):

start by setting the function equal to 10

OpenStudy (sanra123):

10=-16t^2+20t+6

OpenStudy (sanra123):

Like that?

ganeshie8 (ganeshie8):

yes subtract 10 both sides

ganeshie8 (ganeshie8):

and maybe cancel any common constant factors

ganeshie8 (ganeshie8):

you will get a quadratic

ganeshie8 (ganeshie8):

you can solve a quadratic by factoring or by using quadratic formula

OpenStudy (sanra123):

-16t^2+20t=4

ganeshie8 (ganeshie8):

\[10=-16t^2+20t+6\] subtracting 10 both sides gives \[0=-16t^2+20t-4\] right ?

OpenStudy (sanra123):

Yes

ganeshie8 (ganeshie8):

cancel -4 through out, you get \[4t^2-5t+1 = 0\]

OpenStudy (sanra123):

Why do we cancel 4

OpenStudy (sanra123):

Then how do we factor it

ganeshie8 (ganeshie8):

you're allowed to divide a non zero number both sides

ganeshie8 (ganeshie8):

dividing -4 through out gives you small numbers

OpenStudy (sanra123):

Oh ok

ganeshie8 (ganeshie8):

small numbers are easy to handle

OpenStudy (sanra123):

So then how do we factor it

ganeshie8 (ganeshie8):

\[4t^2-5t+1 = 0 \] pick two numbers such that : 1) they add up to -5 2) they multiply to 4

OpenStudy (sanra123):

I mean -1 -4

ganeshie8 (ganeshie8):

yes! write the middle term using them \[4t^2\color{Red}{-5t}+1=0\]

ganeshie8 (ganeshie8):

\[4t^2\color{Red}{-1t - 4t}+1=0\]

ganeshie8 (ganeshie8):

factor the GCF from first two terms and the last two terms

OpenStudy (sanra123):

I never took this method can't. We use quadratic formula

ganeshie8 (ganeshie8):

okay use quadratic formula then

OpenStudy (sanra123):

Do we use the simplest factor

OpenStudy (sanra123):

4t^2-5t+1

ganeshie8 (ganeshie8):

yes

ganeshie8 (ganeshie8):

work it fast, otherwise @Jhannybean going to forcibly teach you completing the square !

OpenStudy (sanra123):

Hahaha ok

OpenStudy (sanra123):

I got 5+-sqrt5 /8

ganeshie8 (ganeshie8):

doesn't look correct, try again

OpenStudy (jhannybean):

\[4t^2-5t+1 = 0\]\[t^2 -5t+4=0\]\[(t-4)(t-1)=0\] Since we multiplied the 4 over to the one to make it a quadratic equation that's easier to solve, we have to divide by the coefficient as well. \[\left(t-\frac{4}{4}\right)\left(t-\frac{1}{4}\right)=0\] simplify the fractions you can, and for those that you cannot, multiply the denominator to the variable. \[(t-1)(4t-1)=0\]

OpenStudy (sanra123):

Soory I didn't understand this method

OpenStudy (sanra123):

:(

OpenStudy (jhannybean):

you're taking a complex quadratic equation and simplifying it in order to find your factoredform

OpenStudy (sanra123):

I still don't learn how to simplify quadratic equations that's why

OpenStudy (jhannybean):

Then I'll teach you this method of quick factorization :)

OpenStudy (sanra123):

Now?

OpenStudy (jhannybean):

Let's work it step by step. you understand how you found the \[4t^2 -5t-1=0\] right?

OpenStudy (sanra123):

No

OpenStudy (jhannybean):

So you did not follow what @ganeshie8 was guiding you through?

OpenStudy (sanra123):

Wait let we go back and check

OpenStudy (jhannybean):

Take a minute to review the stuff up there ^ when you're done looking through it, let me know.

OpenStudy (sanra123):

i did simplify it in another way 4(4t^2+5t-1)=0 4 will be canceled

OpenStudy (sanra123):

then we are left with 4t^2+5t-1

OpenStudy (sanra123):

@Jhannybean

OpenStudy (sanra123):

but then i don't know how to simplify more

OpenStudy (jhannybean):

Okay, do you do understand how to get to this point.

OpenStudy (sanra123):

yess

OpenStudy (jhannybean):

I find it really hard to factor out a quadratic function that has a coefficient in the front, in this case a 4.

OpenStudy (sanra123):

so is there another way

OpenStudy (jhannybean):

So what I do is this simple trick, but you have to constantly remember that you are modifying the equation to make it easier to solve, but the function WILL NOT CHANGE.

OpenStudy (sanra123):

ok

OpenStudy (jhannybean):

This way that I am teaching you is probably the easiest method of going about factoring out a quadratic equation without having to use the quadratic formula

OpenStudy (sanra123):

yay im so happy

OpenStudy (jhannybean):

So in order to make it easier to factor, I simply took the 4 that was my coefficient and multiplied it t the constant. This made it easier to find two numbers that add to give me -5, and multiply to give me +4

OpenStudy (jhannybean):

\[t^2 -5t+4=0\]That is how I got to this step. following me so far?

OpenStudy (sanra123):

1 min

OpenStudy (sanra123):

so we divide the coefficient with 4 and times the last number by 4

OpenStudy (sanra123):

but then why did the signs change

OpenStudy (jhannybean):

In that step we are simply 'creating' an easier quadratic function to factor.Would you agree this: \[4t^2 -5t+1\] is a lot harder to factor than \[t^2-5t+4\]

OpenStudy (jhannybean):

Just follow me step by step, do not jump to conclusions just yet.

OpenStudy (sanra123):

ok so signs didn't change

OpenStudy (jhannybean):

No, they didn't.

OpenStudy (sanra123):

oh ok

OpenStudy (jhannybean):

So now you understand how we came to \(t^2-5t+4\), yes?

OpenStudy (sanra123):

yes

OpenStudy (jhannybean):

Okay.

OpenStudy (jhannybean):

Now we find two numbers that ADD to give -5, but MULTIPLY to give -4. What can those be?

OpenStudy (sanra123):

but wait why -4

OpenStudy (sanra123):

u mean 4

OpenStudy (sanra123):

so its -1 -4

OpenStudy (jhannybean):

yes,typo,sorry. Glad you caught that

OpenStudy (jhannybean):

yes, -4 , -1.

OpenStudy (sanra123):

ok wait before we continue if we dot have a leading coefficient we divide by and mutually by what

OpenStudy (sanra123):

dont*

OpenStudy (jhannybean):

If we don't have a leadng coefficient, then we just factor the quadratic like normal

OpenStudy (jhannybean):

The only reason we are adding on a few extra steps like simplifying the quadratic is just to make it easier to factor.

OpenStudy (jhannybean):

But moving on, we will have \((t-1)(t-4)\) , but if you expand this, you will only have the simplified quadratic function, \(t^2-5t+4\)

OpenStudy (sanra123):

so this method is for if we have leading coefficient

OpenStudy (jhannybean):

Yes! :) There you go.

OpenStudy (sanra123):

:) ok lets continue

OpenStudy (jhannybean):

This is where it might get a little tricky, so pay close attention.

OpenStudy (sanra123):

ok

OpenStudy (jhannybean):

Like I mentioned before, we need to use this simplified form to revert back to our original equation: \(4t^2-5t+1\). Because we initially multiplied the 4 to make it easier to factor, we need to now divide by the same coefficient so we can get back the original function

OpenStudy (jhannybean):

Thats why from \[(t-1)(t-4)\] we will divide both of those values by 4.\[\left(t-\frac{1}{4}\right)\left(t-\frac{4}{4}\right)\]

OpenStudy (sanra123):

ok

OpenStudy (sanra123):

so its 1/4 the ANSWER

OpenStudy (jhannybean):

No not yet.

OpenStudy (sanra123):

OH OK

OpenStudy (jhannybean):

After dividing, we cam simplify the fractions, whichever fractions simplify. \(\frac{4}{4} = 1\) and \(\frac{1}{4}\) stays the same.

OpenStudy (sanra123):

oh ok

OpenStudy (jhannybean):

So we will have \[\left(t-\frac{1}{4}\right)\left(t-1\right)\]